OpenStudy (anonymous):

find the zeros of each functin y=(x-2)(x+4)

OpenStudy (anonymous):

The zeroes are always what makes the function equal to zero.

OpenStudy (shamil98):

Set y = 0 and solve for x.

OpenStudy (whpalmer4):

And in a product, any value that makes one of the product terms = 0 is a zero. \[0 = (x-2)(x+4)\]That's true if \((x-2) = 0\) or if \((x+4) = 0\)

OpenStudy (whpalmer4):

You can use this in reverse to construct a function that has the desired zeros. If you want a function that has zeros of 3 and -7, \[f(x) = a(x-3)(x-(-7)) = a(x-3)(x+7)\]There's your function! The value of \(a\) can be 1, or you can choose a different value if you need to have the function have a specific value at a given point. In the following graph, all of the parabolas have the same zeros (aka roots), but a different value for \(a\).